A simple random sample of 60 items resulted in a sample mean of 91. The population standard deviation is 12.

a. Compute the 95% confidence interval for the population mean (to 1 decimal).

(______,_______)?

b. Assume that the same sample mean was obtained from a sample of 120 items. Provide a 95% confidence interval for the population mean (to 2 decimals).

(_______,_______)?
Answer #1

mean = 91

sd = 12

Z for 95% confidence interval = Z0.025 = 1.96

a) n = 60

confidence interval =(\bar{X} \pm Z_{0.025}*\frac{\sigma}{\sqrt{n}})

                               =(91 \pm 1.96*\frac{12}{\sqrt{60}})

                               =(91 \pm 3.03)

                               = (88 , 94)

b) n = 120

confidence interval =(\bar{X} \pm Z_{0.025}*\frac{\sigma}{\sqrt{n}})

                               =(91 \pm 1.96*\frac{12}{\sqrt{120}})

                               =(91 \pm 2.1)

                               = (88.9 , 93.1)